OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below. Show your work. (3 points) x2 − 16 = 0 and x2 = −2x + 24 Part 2: Describe what the solution(s) represent to the graph of each. (2 points) Part 3: How are the graphs alike? How are they different?

OpenStudy (anonymous):

1) x2 − 16 = 0 (x+4)(x-4)=0 x=4,-4

OpenStudy (anonymous):

Part 1. Solve each quadratic: x² - 16 = 0 ==> x² = 16 ==> x = ±√16 = ±4 ==> x = -4 and x = 4 x² = -2x + 24 ==> x² + 2x - 24 = 0 ==> (x + 6)(x - 4) = 0 ==> x = -6 and x = 4 The solutions (roots) to these quadratic equations are the x-intercepts, , where the curves cross the x-axis, where y = 0. Part 2: The first graph, y = x² - 16, represent an upward-opening parabola with it's vertex at (0,-16) and x-intercepts at (-4,0) and (4,0). The second graph, y + x² = -2x + 24, also represents an upward-opening parabola with it's vertex at (-1,-25) and x-intercepts at (-6,0) and (4,0).

OpenStudy (anonymous):

x2 +2x -24=0 x2 +6x -4x-24=0 (x+6)(x-4)=0 x=4,-6

OpenStudy (anonymous):

thankyou double driv do you know part 3 ?

OpenStudy (anonymous):

both represents paarabola with axis parallel to y-axis and each passes through 4,0

OpenStudy (anonymous):

graph them and compare them

OpenStudy (anonymous):

okay thanks guys